Eight years ago, Heather was twice as old as her brother David. Today she is 13 years older than him. How old is she?
Answer by ptaylor(2198)

Let x=David's current age; x-8=David's age 8 years ago
Then x+13=Heather's current age; x+13-8=x+5=Heather's age 8 years ago
Soooooo
x+5=2(x-8)get rid of parens
x+5=2x-16 subtract 2x and also 5 from each side
-x=-21
x=21----David's current age
x+13=21+13=34----Heather's current age
CK
Today: 21+13=34 ok
8 years ago: 34-8=2(21-8)
26=2*13 ok
Hope this helps----ptaylor